OpenStudy (anonymous):

helppp! please EXPLAIN and do not just give me the answer(: it says find the volume of the sphere use 3.14 for PI. the r= 3 cm

OpenStudy (anonymous):

you want me to derive the formula to find the volume of sphere

OpenStudy (kropot72):

Do you know the formula to find the volume of a sphere?

OpenStudy (anonymous):

no i don't know the formula.

OpenStudy (sw050399):

It's: V = (4/3)(pi)(r^3)

OpenStudy (anonymous):

Think of a sphere as a bunch of disks being stacked onto each other. Each of these disks represents a circle with a radius of |r|<3cm, meaning that the radius of each disk is at most 3cm. So, we what we want to do is sum the area of each individual disk to obtain the volume of the sphere. As to how you do that, let me show you: 1. We know that the area of a circle is 2*pi*r, gimme a sec. OK wait, what level math are you?

OpenStudy (anonymous):

If what you need to do is simply apply the formula, then use what sq050399 gave you, but if you need actually derive this formula and explain how it's used, let me know.

OpenStudy (anonymous):

Using the formula \[V=\frac{4}{3}\pi\cdot{r^3}\] simply plug in r = 3, to obtain the volume.

OpenStudy (kropot72):

The volume of a sphere is \[V(sphere)=\frac{4}{3}\pi r ^{3}\] Just plug in 3 for r and calculate the volume in cubic centimeters.

OpenStudy (anonymous):

is the answer v=113.04?

OpenStudy (anonymous):

Yes, but to be precise, you want to say it's 113.04 cm^3.